A total of
492
tickets were sold for the school play. They were either adult tickets or student tickets. There were
58
fewer student tickets sold than adult tickets. How many adult tickets were sold?

Answer by FrankM(1040)

A=S+58
A+S=492
S+58+S=492
2S+58=492
2S=434
S= 217
A= 275